# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v1

@@ VII Tool Usability
OODTE  can be used directly to test whether the ONNX Optimizer generates optimized models without negatively affecting their accuracy. Specifically, OODTE can verify that the ONNX Optimizer performs correctly on real-scale models of various types and architectures. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type.
OODTE is an automated utility designed to test the ONNX Optimizer. It can be configured to perform tests both manually and automatically. As a configuration-based tool, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added, whether by the optimizer developers or users implementing custom passes.
In terms of models and datasets, OODTE is adaptable to models provided both locally and fetched from online repositories. By default, OODTE loads models from the official ONNX Model Hub [8], which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [7], it can fetch and utilize models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using widely used packages like the datasets [4] package or from local storage. Additionally, OODTE can be easily extended to support any dataset suitable for testing.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and in isolation.
@@ VIII Threats To Validity
In terms of threats to validity, we considered and addressed the following aspects: (1) Comprehension of Experiment Set: We sourced $130$ models from the official ONNX Model Hub, covering classification, object detection, and text generation—categories widely used across various domains. These models are well-established and mature, as the main purpose of the hub is to provide researchers and developers with easily accessible models that require minimal configuration and fine-tuning. The hub offers a broad range of ONNX model versions (opsets), which we considered in our experiments, ranging from version $7$ (the lowest version compatible with the current ONNX Runtime at the time of writing) to version $12$ (the highest available via the ONNX Hub API).
We believe our experiment set is a representative and comprehensive set of models for testing and experimentation. However, OODTE supports models from any hub that complies with ONNX API specification. While our experiment set does not include all models in the hub, we believe it is representative. We excluded models with an opset version lower than $7$ (which are incompatible with the current ONNX and ONNX Runtime version), as well as models for body analysis and super-resolution, since the selected classification and object detection models adequately cover these areas. Additionally, models that only prepare inputs for preprocessing were excluded for being non-generalizable, as were certain transformer models, which were represented by GPT-2.
Overall, we believe our experiment set is comprehensive, representing models commonly used by developers and researchers in the community and industry across a wide range of domains, while adhering to the latest versions supported by ONNX. (2) Optimization Order: We acknowledge that the optimization order can impact model performance [31], a common challenge in compiler research. However, exploring this aspect is beyond the scope of this work. To ensure the validity of our approach, we used two modes in OODTE: (1) the default ONNX Optimizer configuration (which includes the fuse and eliminate passes), and (2) runs where all $47$ optimizer passes were applied atomically. While our approach does not address potential defects arising from the combination of passes, it has been shown to be effective towards detecting correctness issues. (3) Proper Optimization Usage: We followed all guidelines provided by the ONNX Optimizer API to ensure the correct use of each pass. In cases where a pass was expected to produce model graphs that were not strictly compliant with the ONNX format, we took note of this and clearly indicated such instances when reporting the respective issues to the developers. (4) CPU Devices Used: We used both an Intel i5 and an Intel i7 in parallel for our experiments. To ensure consistency across devices, we (a) randomly sampled a subset of models of each type and ran them on both devices, verifying that the results were consistent, and (b) referred to the work of Louloudakis et al. [41], who observed that accuracy remains consistent across different processing devices using optimization. This supported our finding from the sample that results remain consistent across devices.
@@ IX Discussion
We make the following observations from our experiments:
(1) object detection models are more prone to lead into a crash during the optimization process. The vast majority of the crashes we detected were related to object detection and semantic segmentation models. We attribute this to the increased complexity of object detection models in comparison to those utilized only for classification tasks, which are typically simpler in terms of architecture.
(2) The object detection capabilities of problematic models might be considerably affected. In particular:
  • YOLOv3 models show a high average IoU, F1 score, precision, and recall (over 98%). The optimized model misplaces bounding boxes, but not enough to cause misclassifications, as no misclassifications were found in the top-10 label predictions.
  • YOLOv2 showed low IoU values, likely due to its use of multiple bounding boxes per detected object, where small deviations accumulate, negatively affecting the IoU score. However, the high F1 score, average precision, and recall suggest that these deviations do not significantly impact object detection, similarly to how it is illustrated by the left bird instance in Figure 2.
  • In contrast, SSD showed about $4.8$% deviation in F1, precision, and recall between the original and optimized models, leading to $2.9-6$% misclassification rates, despite a high IoU ($>97$%). This is similar to the right bird instance in the illustrative example in Figure 2.
All models presented a considerable mAP score deviation of up to $14.31$%. This indicates that some classes in the model were not properly identified, meaning that the optimizer might affect all object detection and semantic segmentation models towards missing specific objects. This can have a severe effect in models, as despite correctly classifying the objects they detect, they might ”miss” a non-negligible amount objects in the scenes presented to them.
(3) Classification models are more prone to differences in accuracy of the optimized model, compared to the original model. We observed two issues that had an adverse effect towards model classification accuracy. The root causes of such deviations, were attributed to two main issues, (a) one related to implicit, improper change of the ONNX format version for older model opset versions - in which case a large amount of classification models were built using these opsets, and (b) one related to the generation of problematic graph nodes, in relation to convolutional layers (fuse_bn_into_conv pass), a pass that has greater effect in classification models.
@@
(4) Text generation models were affected only to a negligible extent. From the $8$ model instances under test related to text generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model ($4$ when selecting the first token, $2$ when selecting the second and another $2$ when selecting the third token on each prediction). However, we consider this result negligible, as a common practice is to use random token selection across token predictions, and such a difference is so negligible it will essentially not affect model behavior.
Overall, we recognize that the ONNX Optimizer is a tool that performs well for most of the passes and produces good results. We aspire to contribute with this publication, so that the reported issues get fixed by the optimizer developers.
X Future Work
We identify three potential extensions of this work as future research directions:
(1) Execution Time Measurements: This study primarily focused on optimizer correctness, specifically detecting crashes and accuracy faults. However, given the context of model optimization, it is important to also assess how the optimizer impacts model performance, particularly in terms of execution time (for both CPU and GPU deployments) and other factors such as model size. OODTE can be easily extended to support hardware acceleration model execution (either by using the ONNX Runtime API [3], or by integrating a third-party AI compiler that generates target device code, such as Apache TVM [12]) We leave this as a subject for future work, which could provide valuable insights.
(2) Selection of Different Model Types: This work focused on classification, object detection, and text generation models. However, OODTE can be applied to other model types, such as speech recognition, super-resolution, and body recognition. We consider this as a direction for future work.
(3) Examine Optimization Pass Order: OODTE supports optimization with a subset of passes in a defined order. However, our current approach does not account for this, and it can be explored in future contributions.
XI Conclusion

### v2

@@ VII Tool Usability
OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, OODTE loads models from the official ONNX Model Hub [7], which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [6], it can fetch and use models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using packages like the datasets [3] package or from local storage.
Finally, while OODTE applies a differential testing methodology focused on the ONNX Optimizer, this approach is generalizable and can be easily adapted to other AI optimizers or compilers that meet two primary requirements: (1) they provide a runtime for model execution, and (2) they expose an API that includes all the passes applied during model optimization, both as a bundle and as independent passes.
@@ VIII Threats To Validity
We list the external and internal threats to validity in our evaluation below.
VIII-A External Threats
There are three primary external threats to the validity of our findings.
(1) Model Selection: Our evaluation included 130 models sourced from the official ONNX Model Hub, covering widely used categories such as classification, object detection, and text comprehension and generation. These models were selected for their maturity, accessibility, and broad adoption across academic and industrial domains. While we aimed to create a representative and diverse experimental set, the selection is not exhaustive. Models with opset versions below 7 were excluded due to incompatibility with the current ONNX and ONNX Runtime versions, which may limit our findings to more recent models. Additionally, to keep the evaluation tractable, we did not include certain model types—such as diffusion models and reinforcement learning—which may behave differently under optimization. Although our tool supports any ONNX-compliant model hub, the exclusion of these model types presents a potential threat to the generalizability of our results. We plan to address this limitation in future work.
To keep the experiments tractable, we did not evaluate model types such as diffusion models and reinforcement learning in this work; however, we plan to include them in future studies.
(2) Dataset Selection: Our experimental evaluation relies on widely used, real-world benchmark datasets tailored to the specific tasks of the models under test. While this approach ensures relevance and practical applicability—e.g., using ILSVRC [55] for image classification and object detection models like ResNet [22] and GoogleNet [62], and leveraging established HuggingFace datasets [3] for NLP tasks—it may limit generalizability. The use of popular datasets could introduce bias toward well-studied scenarios and may not reflect performance on less common or emerging tasks.
(3) Devices Used: Our experiments were conducted on two hardware platforms: an Intel i5 and an Intel i7 processor. While we verified consistency by (a) randomly sampling models of each type and running them on both devices with consistent results, and (b) referencing prior work by Louloudakis et al. [41], which found that optimization does not affect accuracy across processing devices, this still presents a potential threat to validity. The limited range of hardware may not capture performance or accuracy variations that could arise on other architectures, such as GPUs or edge devices, and may limit the generalizability of our findings.
VIII-B Internal Threats
We identify the following internal threat to validity in our findings.
Optimization Pass Ordering: The order in which optimization passes are applied can significantly affect model performance and behavior, as highlighted in prior work [31]. This is a well-known challenge in compiler research. While this work does not explore the impact of pass ordering in depth, we adopted two evaluation modes in OODTE: (1) the default ONNX Optimizer configuration, which includes commonly used passes such as fuse and eliminate, and (2) a configuration where all 47 available optimization passes were applied atomically. Although this strategy does not account for potential issues arising from specific pass combinations or orderings, it has proven effective in uncovering correctness issues in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
@@ IX Discussion
We summarize below our findings for the different model types in our experiments.
(1) Object detection models crash more frequently when optimized.
Most crashes during optimization occurred in object detection and semantic segmentation models, rather than in classification models. This is likely due to their greater architectural complexity—being larger and structurally more complex—which leads the ONNX Optimizer to apply more extensive transformations, increasing the likelihood of encountering corner cases that trigger crashes related to the ONNX Optimizer passes.
While text generation and comprehension models are also typically large, manual inspection revealed that the ONNX Optimizer applies only a limited set of optimizations to them. Most passes are skipped, as many of the layer patterns targeted by the optimizer (e.g., convolutions)—common in classification and object detection models—are not typically present in transformer-based architectures used for language tasks.
(2) Optimization can reduce the accuracy of object detection models.
While some object detection models remained relatively stable after optimization, others showed notable performance degradation. For example, the YOLOv3 model maintained high detection accuracy with no misclassifications, although it exhibited minor shifts in bounding box positions. YOLOv2 also showed stable performance, with slightly lower Intersection over Union (IoU) due to its use of multiple bounding boxes per object, but still achieved strong detection metrics overall.
In contrast, the SSD model experienced a significant decline in F1 score, precision, and recall, leading to higher misclassification rates despite retaining a high IoU. Across all models, we observed substantial drops in mean Average Precision (mAP), with reductions of up to 14.31%, suggesting that optimization may cause certain object classes to be missed. This raises concerns about the reliability of optimized models in object detection and segmentation tasks.
(3) Optimizations can affect label accuracy in classification models. We identified two primary causes for differences in label accuracy following optimization. (a) The first relates to the implicit and improper upgrade of the ONNX format version in models using older opset versions. A significant number of classification models were built on these legacy opsets, making them particularly susceptible to this issue. (b) The second involves the generation of problematic graph nodes, particularly in relation to the fuse_bn_into_conv pass. This pass, which fuses batch normalization into convolutional layers, has a strong impact on classification models and can lead to unintended changes in model behavior after optimization.
@@
(4) Outputs from Text comprehension and generation models were not affected by optimization. From the $8$ model instances under test related to text comprehension and generation, we only found one crash issue related to GPT-2, as well as a very small number of samples (8) producing a BLEU score that indicates mismatch between the original and the optimized model, for the top-3 token selections of the model.
Overall, we find ONNX optimizations can result in output differences between the optimized and unoptimized model, with some model types being more susceptible to optimization errors than others. OODTE  was able to successfully isolate the optimization passes responsible for the output differences for different model types.
X Conclusion